# LLAMAFUZZ: Large Language Model Enhanced Greybox Fuzzing

### v3

@@ Abstract
LLAMAFUZZ: Large Language Model Enhanced Greybox Fuzzing
Hongxiang Zhang  and Yuyang Rong  and Yifeng He  and Hao Chen University of California, Davis Davis, California, USA {hxxzhang, PtrRong, yfhe, chen}@ucdavis.edu
1 Introduction
@@
Fuzz testing, also known as fuzzing, is an automated software testing technique that generates test seeds to discover vulnerabilities in the target software applications. In recent years, greybox fuzzing has drawn much attention because of its effectiveness in discovering new vulnerabilities in many programs. As software systems continue to grow in complexity and evolve at an accelerated pace, the need for adapted test inputs has become increasingly important. While bit-level randomized mutation Zalewski (2016); Fioraldi et al. (2020b); Swiecki ; Lyu et al. (2019); Lemieux and Sen (2018) has achieved a lot, they reached a bottleneck in which traditional greybox fuzzers struggle to effectively generate structured data.
@@
General-purpose greybox fuzzers employ high-throughput bit-level mutation. AFL++ Fioraldi et al. (2020b), one of the state-of-the-art greybox fuzzers, combines multiple mutation strategies and scheduling strategies, leading fuzzing to a new level. However, when fuzzing applications that require structured input, blind random bit-level mutation can be problematic. Such mutations often disrupt the integrity of data formats, resulting in inefficient exploration of the input space. As a result, converging to a high and stable coverage and reaching bugs is time-consuming. To accelerate this process, honggfuzz Swiecki shares the file corpus across multi-process and multi-thread execution to boost throughput. However, the effectiveness of merely increasing throughput and adding new random mutations is limited, since the bottleneck is caused by the complex constraints when handling structured seeds. AFL++ and honggfuzz require excessive attempts to generate valid structured inputs.
Therefore, structural awareness of the test seeds is the key to success in fuzzing such software. To generate structured binary seeds, specialized fuzzers use predefined grammars. Gramatron Srivastava and Payer (2021) restructures the grammar to enable unbiased sampling from the input state space and permits more aggressive mutation operations. Moreover, it combines search-based testing to co-evolve the test case generation. However, it requires additional specifications in pre-defined Chomsky Normal Form Chomsky (1959) and Greibach Normal Form Greibach (1965) to construct grammar automata. NAUTILUS Aschermann et al. (2019) employs grammar-aware mutation operators to probe deep program paths and reveal complex bugs, while GRIMOIRE Srivastava and Payer (2021) synthesizes input structures to discover bugs in grammar-based formats.
@@
Figure 1: The overview of fuzzing with LLAMAFUZZ. Each fine-tuning sample is a pair ($S_{i}$, $S_{i}$$^{\prime}$), representing a seed before and after successful mutation. LLAMAFUZZ applies dual-layered, asynchronous mutations—traditional fuzzing and LLM-enhanced. The four light-blue boxes on the right (Execution, Behavior Monitoring, Mutation, Seeds Queue) depict the fuzzing loop.
@@
We integrate LLMs and fuzzer with a lightweight asynchronous job queue, allowing LLAMAFUZZ to run efficiently on single or multiple GPUs. To assess bug-finding capabilities, we compared LLAMAFUZZ against state-of-the-art fuzzers on bug-based benchmark Magma Hazimeh et al. (2020), including AFL++, Moptafl, Honggfuzz, and Fairfuzz. LLAMAFUZZ outperformed its top competitors, discovering 41 bugs on average and 47 unique bugs overall. In addition, we evaluated LLAMAFUZZ on real-world programs across various structured data formats to access its versatility Metzman et al. (2021). LLAMAFUZZ performs competitively with grammar-based fuzzers and outperforms AFL++ on 11 out of 15 fuzzing targets.
Table 1: Mutation example on XML. Format corruptions are highlighted in red, while successful and structurally valid mutations are highlighted in green.
@@
2 Background and related work
Greybox coverage-guided fuzzing tools Zalewski (2016); Fioraldi et al. (2020b); Xu et al. (2024); Pham et al. (2019); Swiecki employ mutation-based strategies to explore input spaces; however, their reliance on random bit-level mutations often leads to inefficiencies and difficulties in generating valid inputs. To mitigate this, these tools use bitmaps to record execution paths and guide mutations toward unexplored code Wang et al. (2019a), although traditional bit-level approaches still struggle with highly structured data. Grammar-based fuzzing Srivastava and Payer (2021); Blazytko et al. (2019); Fioraldi et al. (2020a); Aschermann et al. (2019) overcomes these challenges by using human-specified grammars and operators to generate syntactically valid and diverse inputs.
Recent advances with LLMs offer a promising alternative by effectively capturing complex data structures to guide seed mutations Deng et al. (2023); Xia et al. (2024, 2023); Huang et al. (2024); Yang et al. (2024b), as demonstrated by CHATFUZZ Hu et al. (2023) and compressed-language models Pérez et al. (2024), thereby enhancing both efficiency and bug discovery in fuzzing. Our method leverages LLMs to learn valid input structures, enabling dynamic, structure-preserving mutations. This approach boosts mutation efficiency, code coverage, and bug detection compared to grey-box and grammar-based techniques.
3 Method
3.1 Architecture
@@
3.2 Fine-tune preparation
We followed the LLMs training process by generative pre-training of a language model on a diverse corpus of unlabeled text, followed by discriminative fine-tuning on specific tasks LeCun et al. (2015). The fine-tuning data were collected from real-world fuzzing processes Metzman et al. (2021). We used these data to teach LLM the pattern and mutation of structured data, enabling LLM to modify a given seed to generate valuable seeds while keeping the original structure.
@@ 3.2 Fine-tune preparation
We expected the LLM to be able to understand the structure of data and generate structured seeds for testing, thus we needed to collect a training set first. Specifically, we collected valuable seeds from FuzzBench Metzman et al. (2021) history experiment data and AFL++ fuzzing data that (1) found new paths, (2) had different hit-counts, (3) triggered crashes. The reasons are intuitive: improving coverage will help fuzzer explore target programs to find vulnerabilities in the unvisited path since bugs can not be found in undiscovered paths. While seeds with different hit counts111number of times a seed exercises a specific path may not directly improve coverage, they execute the program in varied ways, potentially uncovering vulnerabilities in already visited paths. Ultimately, the goal of fuzzing is to find vulnerabilities, so any seed that triggers crashes is valuable.
Figure 2: The workflow of dataset pre-processing. Each binary seed pair consists of the original seed $S_{1}$ and its mutated version $S_{1}$$^{\prime}$. Pairs 11 to $n$ on the right represent entries in the fine-tuning dataset.
@@
We construct a generic seed mutation model by converting binary input files to a uniform hexadecimal representation following Pérez et al. (2024). This conversion serves three purposes. First, to enable LLAMAFUZZ to handle various data formats, we need a uniform data reading method, as adapting to each format is impractical. Second, traditional fuzzers operate at the bit level on binary seeds, whereas LLMs typically require natural language input. Therefore, it is essential to convert the training data to a format that LLMs can understand. Third, the conversion is expected to be efficient, as delay would directly impact fuzzing throughput. Compared to other encoding schemes like base64, hexadecimal is more intuitive, easier to convert from binary. Note that this conversion applies only to binary data; for text-based data, we only add prompts to the seeds, as CHATFUZZ Hu et al. (2023) has shown LLM can process them directly.
@@
As aforementioned, each pair of contiguous hexadecimal characters was compiled into a phrase. However, not all such phrases were present in the tokenizer’s vocabulary. To tackle this, we employed a tokenizer that uses the Byte-Pair Encoding algorithm Sennrich et al. (2015); Touvron et al. (2023), which splits unknown phrases into smaller subword units. This allows the LLAMAFUZZ to generate and understand various data formats.
3.3 Fine-tuning LLM for mutation
Fine-tuning pre-trained models is a common paradigm for achieving proficiency in specific downstream tasks LeCun et al. (2015); von Werra et al. (2020); Ding et al. (2023); Han et al. (2024). This process builds upon the pre-trained model’s general understanding and adapts it to specific tasks through supervised learning. Similarly, supervised fine-tuning (SFT) is necessary when employing a general-purpose LLM for structured data mutation.
As shown in Figure 2, Pair $1$ to Pair $n$ provides example prompts to guide the model in structured data mutation. Each prompt consists of the original structured data and its desired mutated result in hexadecimal representation. Subsequently, we prompt the format keywords allowing LLM to take the general understanding of the format from pre-trained knowledge to do the mutation. Prompt examples are provided in AppendixA.3
Following prior works Xia et al. (2024); Yang et al. (2024c), we apply SFT von Werra et al. (2020) to adapt the LLM. The supervised fine-tuning objective is defined as:
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log P\Big(y^{(i)}\mid x^{(i)};\theta\Big)$ (1)
@@ 3.3 Fine-tuning LLM for mutation
Additionally, we apply model quantization Polino et al. (2018) with mixed-precision $fp16$ and integrate LoRA Hu et al. (2021) to speedup training and inference while maintaining accuracy. Training details are provided in AppendixA.1.
3.4 Integrate Fuzzer and LLM
Speed is paramount for greybox fuzzers, which can execute hundreds or even thousands of seeds per second Yun et al. (2018); Zheng et al. (2019). Any additional processes integrated into the greybox fuzzer, could potentially impair overall throughput and negatively impact fuzzing performance. Particularly, LLM generation is slower and more resource-intensive, primarily requiring substantial GPU resources.
To address this speed mismatch, we designed an asynchronous job [^2] Details asynchronous queue design are provided in AppendixA.2 queue for communication between the LLM and the fuzzer. The fuzzing process is formulated as follows:
$Q_{\mathrm{LLM}}$ $=\{(x^{(i)},t_{i})\mid x^{(i)}\in X,\,t_{i}\in T\},$ (2)
@@ 3.4 Integrate Fuzzer and LLM
$x^{i+1}=\operatorname*{\arg\max}_{x\in Q_{\text{LLM}}\cup Q_{\text{AFL}}}\,\{\text{coverage}(x)\}$ (3)
@@
This asynchronous process eliminates waiting time, enabling the fuzzer to run at high speed without delays from LLM mutation tasks. This ensures that integrating the LLM enhances the fuzzer’s capabilities without compromising efficiency. Additionally, the dual-layer structure allows easy replacement of different LLMs. An ablation study is provided in AppendixB.2.
4 Experiment
4.1 Benchmarks and Evaluation Metrics
Magma V1.2
@@ 4.1 Benchmarks and Evaluation Metrics
Hazimeh et al. (2020) is a ground-truth fuzzing benchmark suite based on real programs with real bugs. Table 5 outlines details of fuzz targets, where four columns indicate benchmark, project, fuzz target, and expected file format. In the Magma experiment, we compared LLAMAFUZZ with AFL++, Moptafl, Honggfuzz, and Fairfuzz. All baseline fuzzers except for AFL++ [^4] We used a more recent version of AFL++ (version 61e27c6) than the one provided in Magma, ensuring that we have access to the latest enhancements. were provided in Magma. Following Magma setup, we use the number of detected bugs and the time to trigger them as key metrics.
OSS-Fuzz
@@
We evaluated a set of real-world programs from OSS-Fuzz Serebryany (2017). For fairness and consistency in the evaluation process Klees et al. (2018), we evaluated in a standard benchmark, FuzzBench Metzman et al. (2021). The specific applications are detailed in Table 5. The chosen benchmark encompassed 12 open-source programs that process different structured data in their latest versions. As suggested by Klees et al. (2018) and Hazimeh et al. (2020), each experiment were repeated ten times, each trial lasted for 24 hours. We evaluated code coverage using established measures Böhme et al. (2022); Klees et al. (2018); Wei et al. (2022), ensuring consistency by reporting branch coverage via afl-cov. Statistical significance was analyzed using the Mann-Whitney U test Klees et al. (2018), and the Vargha-Delaney statistic Olsthoorn et al. (2020) quantified effect size. Further details are provided in AppendixA.5.
4.2 Implementation details

### v4

@@ Abstract.
Greybox Fuzzing, Large Language Models
copyright: acmlicensedjournalyear: 2026copyright: ccconference: 7th ACM/IEEE International Conference on Automation of Software Test (AST 2026); April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 7th ACM/IEEE International Conference on Automation of Software Test (AST 2026) (AST ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793654.3793756isbn: 979-8-4007-2476-3/2026/04ccs: Security and privacy Software security engineeringccs: Software and its engineeringccs: Computing methodologies Natural language processing
1. Introduction
@@
Fuzz testing, also known as fuzzing, is an automated software testing technique that generates test seeds to discover vulnerabilities in the target software applications. In recent years, greybox fuzzing has drawn much attention because of its effectiveness in discovering new vulnerabilities in many programs. As software systems continue to grow in complexity and evolve at an accelerated pace, the need for adapted test inputs has become increasingly important. While bit-level randomized mutation (Zalewski, 2016; Fioraldi et al., 2020b; Swiecki, ; Lyu et al., 2019; Lemieux and Sen, 2018) has achieved a lot, they reached a bottleneck in which traditional greybox fuzzers struggle to effectively generate structured data.
@@
General-purpose greybox fuzzers employ high-throughput bit-level mutation. AFL++ (Fioraldi et al., 2020b), one of the state-of-the-art greybox fuzzers, combines multiple mutation strategies and scheduling strategies, leading fuzzing to a new level. However, when fuzzing applications that require structured input, blind random bit-level mutation can be problematic. Such mutations often disrupt the integrity of data formats, resulting in inefficient exploration of the input space. As a result, converging to a high and stable coverage and reaching bugs is time-consuming. To accelerate this process, honggfuzz (Swiecki, ) shares the file corpus across multi-process and multi-thread execution to boost throughput. However, the effectiveness of merely increasing throughput and adding new random mutations is limited, since the bottleneck is caused by the complex constraints when handling structured seeds. AFL++ and honggfuzz require excessive attempts to generate valid structured inputs. Therefore, structural awareness of the test seeds is the key to success in fuzzing such software. To generate structured binary seeds, specialized fuzzers use predefined grammars. Gramatron (Srivastava and Payer, 2021) restructures the grammar to enable unbiased sampling from the input state space and permits more aggressive mutation operations. Moreover, it combines search-based testing to co-evolve the test case generation. However, it requires additional specifications in pre-defined Chomsky Normal Form (Chomsky, 1959) and Greibach Normal Form (Greibach, 1965) to construct grammar automata. NAUTILUS (Aschermann et al., 2019) employs grammar-aware mutation operators to probe deep program paths and reveal complex bugs, while GRIMOIRE (Srivastava and Payer, 2021) synthesizes input structures to discover bugs in grammar-based formats.
@@
Figure 1. The overview of fuzzing with LLAMAFUZZ. Each fine-tuning sample is a pair ($S_{i}$, $S_{i}$$^{\prime}$), representing a seed before and after successful mutation. LLAMAFUZZ applies dual-layered, asynchronous mutations—traditional fuzzing and LLM-enhanced. The four light-blue boxes on the right (Execution, Behavior Monitoring, Mutation, Seeds Queue) depict the fuzzing loop.
The overview of fuzzing with LLAMAFUZZ. Each fine-tuning sample is a pair ($S_{i}$$^{\prime}$\$), representing a seed before and after successful mutation. LLAMAFUZZ applies dual-layered, asynchronous mutations—traditional fuzzing and LLM-enhanced. The four light-blue boxes on the right (Execution, Behavior Monitoring, Mutation, Seeds Queue) depict the fuzzing loop.
@@
We integrate LLMs and fuzzer with a lightweight asynchronous job queue, allowing LLAMAFUZZ to run efficiently on single or multiple GPUs. To assess bug-finding capabilities, we compared LLAMAFUZZ against state-of-the-art fuzzers on bug-based benchmark Magma (Hazimeh et al., 2020), including AFL++, Moptafl, Honggfuzz, and Fairfuzz. LLAMAFUZZ outperformed its top competitors, discovering 41 bugs on average and 47 unique bugs overall. In addition, we evaluated LLAMAFUZZ on real-world programs across various structured data formats to access its versatility (Metzman et al., 2021). LLAMAFUZZ performs competitively with grammar-based fuzzers and outperforms AFL++ on 11 out of 15 fuzzing targets.
Table 1. Mutation example on XML. Format corruptions are highlighted in red, while successful and structurally valid mutations are highlighted in green.
@@
2. Background
Greybox coverage-guided fuzzing tools (Zalewski, 2016; Fioraldi et al., 2020b; Xu et al., 2024; Pham et al., 2019; Swiecki, ) employ mutation-based strategies to explore input spaces; however, their reliance on random bit-level mutations often leads to inefficiencies and difficulties in generating valid inputs. To mitigate this, these tools use bitmaps to record execution paths and guide mutations toward unexplored code (Wang et al., 2019a), although traditional bit-level approaches still struggle with highly structured data. Grammar-based fuzzing (Srivastava and Payer, 2021; Blazytko et al., 2019; Fioraldi et al., 2020a; Aschermann et al., 2019) overcomes these challenges by using human-specified grammars and operators to generate syntactically valid and diverse inputs.
Recent advances with LLMs offer a promising alternative by effectively capturing complex data structures to guide seed mutations (Deng et al., 2023; Xia et al., 2024, 2023; Huang et al., 2024; Yang et al., 2024b), as demonstrated by CHATFUZZ (Hu et al., 2023) and compressed-language models (Pérez et al., 2024), thereby enhancing both efficiency and bug discovery in fuzzing. Our method leverages LLMs to learn valid input structures, enabling dynamic, structure-preserving mutations. This approach boosts mutation efficiency, code coverage, and bug detection compared to grey-box and grammar-based techniques.
3. Method
3.1. Architecture
@@
3.2. Fine-tune preparation
We followed the LLMs training process by generative pre-training of a language model on a diverse corpus of unlabeled text, followed by discriminative fine-tuning on specific tasks (LeCun et al., 2015). The fine-tuning data were collected from real-world fuzzing processes (Metzman et al., 2021). We used these data to teach LLM the pattern and mutation of structured data, enabling LLM to modify a given seed to generate valuable seeds while keeping the original structure.
@@ 3.2. Fine-tune preparation
We expected the LLM to be able to understand the structure of data and generate structured seeds for testing, thus we needed to collect a training set first. Specifically, we collected valuable seeds from FuzzBench (Metzman et al., 2021) history experiment data and AFL++ fuzzing data that (1) found new paths, (2) had different hit-counts, (3) triggered crashes. The reasons are intuitive: improving coverage will help fuzzer explore target programs to find vulnerabilities in the unvisited path since bugs can not be found in undiscovered paths. While seeds with different hit counts111number of times a seed exercises a specific path may not directly improve coverage, they execute the program in varied ways, potentially uncovering vulnerabilities in already visited paths. Ultimately, the goal of fuzzing is to find vulnerabilities, so any seed that triggers crashes is valuable.
Figure 2. The workflow of dataset pre-processing. Each binary seed pair consists of the original seed $S_{1}$ and its mutated version $S_{1}$$^{\prime}$. Pairs 11 to $n$ on the right represent entries in the fine-tuning dataset.
[…]
@@
We construct a generic seed mutation model by converting binary input files to a uniform hexadecimal representation following (Pérez et al., 2024). This conversion serves three purposes. First, to enable LLAMAFUZZ to handle various data formats, we need a uniform data reading method, as adapting to each format is impractical. Second, traditional fuzzers operate at the bit level on binary seeds, whereas LLMs typically require natural language input. Therefore, it is essential to convert the training data to a format that LLMs can understand. Third, the conversion is expected to be efficient, as delay would directly impact fuzzing throughput. Compared to other encoding schemes like base64, hexadecimal is more intuitive, easier to convert from binary. Note that this conversion applies only to binary data; for text-based data, we only add prompts to the seeds, as CHATFUZZ (Hu et al., 2023) has shown LLM can process them directly.
@@
As aforementioned, each pair of contiguous hexadecimal characters was compiled into a phrase. However, not all such phrases were present in the tokenizer’s vocabulary. To tackle this, we employed a tokenizer that uses the Byte-Pair Encoding algorithm (Sennrich et al., 2015; Touvron et al., 2023), which splits unknown phrases into smaller subword units. This allows the LLAMAFUZZ to generate and understand various data formats.
3.3. Fine-tuning LLM for mutation
Fine-tuning pre-trained models is a common paradigm for achieving proficiency in specific downstream tasks (LeCun et al., 2015; von Werra et al., 2020; Ding et al., 2023; Han et al., 2024). This process builds upon the pre-trained model’s general understanding and adapts it to specific tasks through supervised learning. Similarly, supervised fine-tuning (SFT) is necessary when employing a general-purpose LLM for structured data mutation.
As shown in Figure 2, Pair $1$ to Pair $n$ provides example prompts to guide the model in structured data mutation. Each prompt consists of the original structured data and its desired mutated result in hexadecimal representation. Subsequently, we prompt the format keywords allowing LLM to take the general understanding of the format from pre-trained knowledge to do the mutation.
Following prior works (Xia et al., 2024; Yang et al., 2024c), we apply SFT (von Werra et al., 2020) to adapt the LLM. The supervised fine-tuning objective is defined as:
(1) $\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log P\Big(y^{(i)}\mid x^{(i)};\theta\Big)$
@@ 3.3. Fine-tuning LLM for mutation
Additionally, we apply model quantization (Polino et al., 2018) with mixed-precision $fp16$ and integrate LoRA (Hu et al., 2021) to speedup training and inference while maintaining accuracy.
3.4. Integrate Fuzzer and LLM
Speed is paramount for greybox fuzzers, which can execute hundreds or even thousands of seeds per second (Yun et al., 2018; Zheng et al., 2019). Any additional processes integrated into the greybox fuzzer, could potentially impair overall throughput and negatively impact fuzzing performance. Particularly, LLM generation is slower and more resource-intensive, primarily requiring substantial GPU resources.
To address this speed mismatch, we designed an asynchronous job queue for communication between the LLM and the fuzzer. The fuzzing process is formulated as follows:
(2) $Q_{\mathrm{LLM}}$ $=\{(x^{(i)},t_{i})\mid x^{(i)}\in X,\,t_{i}\in T\},$
@@ 3.4. Integrate Fuzzer and LLM
(3) $x^{i+1}=\max_{x\in Q_{\text{LLM}}\cup Q_{\text{AFL}}}\,\{\text{coverage}(x)\}$
@@
This asynchronous process eliminates waiting time, enabling the fuzzer to run at high speed without delays from LLM mutation tasks. This ensures that integrating the LLM enhances the fuzzer’s capabilities without compromising efficiency. Additionally, the dual-layer structure allows easy replacement of different LLMs.
3.5. Asynchronous Queue Design
As summarized in Algorithm 1, our hybrid fuzzer orchestrates three concurrent components: Main Fuzzing Loop, LLM Mutation Thread, and AFL Mutation Thread. This design decouples resource‐intensive, semantics‐driven mutations from the high‐throughput fuzzing loop, enabling non‐blocking operation and easy scaling or replacement of the LLM backend.
Main Fuzzing Loop.
LLM Mutation Thread.
AFL Mutation Thread.
Input: Initial seed corpus
Output: Discovered interesting test cases
Initialization: start_fuzzer(), llm_mutate_thread(), afl_mutate_thread()
start_fuzzer() while true do
 if Queue_LLM not empty then
      seed $\leftarrow$ dequeue(Queue_LLM)
 else
      seed $\leftarrow$ dequeue(Queue_AFL)
 if run_fuzzer(seed) is interesting then
      enqueue(Queue_LLM_unmutated, seed);
      enqueue(Queue_AFL_unmutated, seed);
llm_mutate_thread() while true do
 if Queue_LLM_unmutated not empty then
      seed $\leftarrow$ dequeue(Queue_LLM_unmutated);
      hex_seed $\leftarrow$ convert_to_hex(seed);
      mutated_seed $\leftarrow$ LLM_generate(hex_seed);
    // GPU
      enqueue(Queue_LLM, mutated_seed);
afl_mutate_thread() while true do
 if Queue_AFL_unmutated not empty then
      seed $\leftarrow$ dequeue(Queue_AFL_unmutated);
      mutated_seed $\leftarrow$ AFL_generate(seed);
    // CPU
      enqueue(Queue_AFL, mutated_seed);
Algorithm 1 Hybrid Fuzzer with LLM- and AFL-based Mutations
4. Experiment
4.1. Benchmarks and Evaluation Metrics
Magma V1.2.
@@ 4.1. Benchmarks and Evaluation Metrics
(Hazimeh et al., 2020) is a ground-truth fuzzing benchmark suite based on real programs with real bugs. Table 2 outlines details of fuzz targets, where four columns indicate benchmark, project, fuzz target, and expected file format. In the Magma experiment, we compared LLAMAFUZZ with AFL++, Moptafl, Honggfuzz, and Fairfuzz. All baseline fuzzers except for AFL++ [^3] We used a more recent version of AFL++ (version 61e27c6) than the one provided in Magma, ensuring that we have access to the latest enhancements. were provided in Magma. Following Magma setup, we use the number of detected bugs and the time to trigger them as key metrics.
OSS-Fuzz.
@@
We evaluated a set of real-world programs from OSS-Fuzz (Serebryany, 2017). For fairness and consistency in the evaluation process (Klees et al., 2018), we evaluated in a standard benchmark, FuzzBench (Metzman et al., 2021). The specific applications are detailed in Table 2. The chosen benchmark encompassed 12 open-source programs that process different structured data in their latest versions. As suggested by Klees et al. (2018) and Hazimeh et al. (2020), each experiment were repeated ten times, each trial lasted for 24 hours. We evaluated code coverage using established measures (Böhme et al., 2022; Klees et al., 2018; Wei et al., 2022), ensuring consistency by reporting branch coverage via afl-cov. Statistical significance was analyzed using the Mann-Whitney U test (Klees et al., 2018), and the Vargha-Delaney statistic (Olsthoorn et al., 2020) quantified effect size.
Experiment Setup and Metrics.
Table 2. Targets information. The programs tested under Magma are utilized in their default versions. For programs included in the real-world benchmark, we specify the exact versions used by listing the Git SHA identifiers.
Project & version Fuzz Target File format
Magma V1.2 libpng libpng_read_fuzzer PNG
libsndfile sndfile_fuzzer Audio
libtiff tiff_read_rgba_fuzzer TIFF
tiffcp
libxml2 read_memory_fuzzer XML
xmllint
lua targets lua Lua
openssl asn1, asn1parse, bignum, Binary blobs
server, client, x509
php json JSON
exif EXIF
unserialize Serialized object
parser PHP
poppler pdf_fuzzer, pdfimages PDF
pdftoppm
sqlite3 sqlite3_fuzz SQL query
Real-world programs mruby 14c2179 mruby_fuzzer mruby
php afa034d php_exec PHP
quickjs 91459fb eval JavaScript
binutils 7320840 fuzz_nm, fuzz_objcopy ELF, String
fuzz_readelf, fuzz_strings
bloaty 34f4a66 fuzz_target ELF, Mach-O
freetype2 cd02d35 ftfuzzer TTF, OTF, WOFF
grok b9286c2 decompress_fuzzer JPEG 2000
kamailio 3f774f3 fuzz_parse_msg SIP message
fuzz_uri URI
libavc 828cdb7 avc_dec_fuzzer AVC
mvc_dec_fuzzer MVC
svc_dec_fuzzer SVC
libhevc d0897de hevc_dec_fuzzer HEVC
openh264 1c23887 decoder_fuzzer H.264 / MPEG-4 AVC
zlib 0f51fb4 uncompress_fuzzer Zlib compressed
We chose Magma for several reasons. First, Magma involves a wide range of popular programs that process diverse structured data with real-world environments, including 9 libraries and 21 objects. Second, unlike LAVA-M (Dolan-Gavitt et al., 2016), which primarily employs synthetic bugs and magic byte comparisons, Magma offers a diverse range of real vulnerabilities, with a total of 138 bugs spanning integer errors, divide-by-zero faults, memory overflows, use-after-free, double-free, and null-pointer dereference scenarios. It incorporates real-world bugs from older versions of software updated to their latest releases, ensuring the benchmark’s relevance and practical applicability. Third, Magma focuses on bug counts and time-to-bug metrics as more direct surrogates for performance (Hazimeh et al., 2020). In addition, we chose FuzzBench in later real-world experiments, because it employed Docker containers to standardize the testing environment for each fuzzer. This setup guaranteed fairness that all fuzzers operated under identical conditions, thus ensuring comparability of results.
To assess the statistical significance of our results, we applied the Mann-Whitney U Test (Klees et al., 2018). As per the Mann-Whitney U-test, a result is statistically significant if the p-value is less than 0.05. In addition, we used the Vargha-Delaney statistic (Olsthoorn et al., 2020) to quantify effect size, providing insight into the magnitude of observed differences. The result was classified as negligible if $\hat{A}_{12}$ was less than 0.5, as medium if it was greater than 0.5 but not exceeding 0.8, and as large if it was greater than 0.8.
4.2. Implementation details